HUBBLE SPACE TELESCOPE DAILY REPORT #2319 PERIOD COVERED: 0000Z (UTC) 02/01/99 - 0000Z (UTC) 03/02/99 Daily Status Report as of 061/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6461 (High-Resolution Next Bright Galactic Nova) The WF/PC-2 was used to make high resolution images of the bright galactic nova U-Scorpio after it erupted to search for fluorescence and reflection light echoes from, and small-scale structure in, the nova circumstellar gas and dust ejecta. These observations were a target of opportunity, and a new SMS had to be built. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7337 (The Central Mass Profile in the Lowest Redshift Cluster Lenses) The WF/PC-2 was used to make observations of the X-ray luminous cluster of galaxies around the radio source PKS 0745-191. This cluster produces the brightest gravitational arc known. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.5 Completed WF/PC-2 6812 (The Nature of Very Red Field Galaxies) The WF/PC-2 was used to make observations of the galactic -cluster 0732. A near-infrared selection is ideal for the study of distant galaxy evolution because it minimizes bias toward finding strongly evolving or star forming galaxies. The observations were completed as planned, and no anomalies were reported. 1.6 Completed STIS/MA2 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD/MA2 7657 (STIS Spectroscopic and Imaging Sensitivity, MAMA) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make observations of GD71 in order to study the spectroscopic and imaging sensitivity. Sensitivity measurements are done for all primary tilts of the gratings and the prism used with the MAMAs, except for the G140L and G230L gratings. Imaging sensitivity measurements are done with all apertures, including a neutral density filter. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: The SSR EDAC error counter was cleared once per ROP SR-1. The Gyro-3 drift rate bias trend shows an increase since the last report. The rate of change over the last two days is 29.16 arcsec/hour/day (was 24.75 arcsec/hour/day). 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH